OpenStudy (anonymous):

solve: 2xyy'=y^2 +xy

OpenStudy (australopithecus):

that isn't right this is calculus

OpenStudy (australopithecus):

or implicit differentiation

OpenStudy (australopithecus):

Yes implicit differentiation is used to find tangents on equations that are not functions (fail vertical line test) It has been so long since I have done such problems it is hard for me to remember how to go about them

OpenStudy (australopithecus):

actually this may not be implicit differentiation.

OpenStudy (unklerhaukus):

The equation can be arragnge to the form dy/dx = f(y/x) apply the subsitution y/x = v y = vx dy/dx = v'x+v v'x+v = f(v)

OpenStudy (unklerhaukus):

this will now be variables seperable